OpenStudy (anonymous):

some oranges are shared among 5 people each person gets three more than he or she would have gotten if the oranges were shared among 8 people how many oranges are there

OpenStudy (anonymous):

Let \(x\) be the number of oranges.

OpenStudy (anonymous):

\[ \frac x5+3 =\frac x8 \]

OpenStudy (anonymous):

thanks i think i figured it out